Grade 12Physical Chemistry

what trasition in the hydrogen spectrum would have same wavelength as a Balmer transition n1 = 4 , n2 = 2, He+ spectrum?

sima roy
8 Years agoGrade 12

1 Answer

Ravneet
8 Years ago
 For He+ ion,
1/λ=z2 [ 1/n1-1/n2 ] R                  WHERE R IS RYDBERG CONSTANT = 109678 m-1
                                                                      Z IS ATOMIC NUMBER
1/λ  =  22*109678[1/22-1/42]
1/λ =   3/4*109678   ----------------------(1)
 
for H atom, (Z = 1)
1/λ= 109678[ 1/n1-1/n2 ]   ----------------------------------  using equation 1 
1/λ= 3/4*109678 = 109678[ 1/n1-1/n2 ]  
now, 
¾ =  1/n1-1/n2 
just do hit and trial , n= 1   and  n2  = 2
or if option is available then check by options.